### v1

@@ Methods
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p^{\theta}(x_{1}|x_{t},t)\right],$ (4)
where $x_{1}$ refers to clean (un-noised) data, $p(t)$ is the noise schedule for $t\in[0,1]$, and $x_{t}$ the noised data. Campbell et al. [38] show that when the forward noising process is applied independently across each dimension of $x_{1}$, the denoising model, $p^{\theta}(x_{1}|x_{t},t)$, can be learned as a factorized over each dimension, $i\in\{1,2,...,D\}$, of the data:
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t},t)\right].$ (5)
Our goal is to demonstrate that this loss function is equivalent to the loss function used by a class of masked language models (MLMs), that we will refer to as “generative masked language models”. The key distinction between generative masked langauge models and their non-generative counterparts, is that during training, generative MLMs mask tokens with a masking probability, $t\in[0,1]$, sampled from a distribution $p(t)$ that has full support on the interval $[0,1]$. This requirement is met by MLMs such as ESM3, but, importantly, is not met by BERT-style MLMs, such as ESM2 as these models use a fixed percentage of tokens that are masked at each step. Letting, $x_{t}$ refer to the masked sequence, we have that the generative masked language model objective function is
$\mathcal{L}_{\text{MLM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t})\right],$ (6)
@@
For any position $i\in U_{t}$, we know with certainty that $x_{1}^{i}=x_{t}^{i}$ by the definition of our forward process. The key observation is that the time variable $t$ only influences which tokens are masked (the masking pattern), but once we observe $x_{t}$, we have complete information about which positions are masked and which are not. The time variable, therefore, provides no additional information needed to predict the clean data. Using the mathematical argument from Gat et al. [39], we can prove this as follows:
@@
$=\frac{\cancel{\big{[}\prod_{i\in M_{t}}(1-t)\prod_{i\in U_{t}}t\big{]}}\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=x_{1}^{i})\big{]}p(t)p(x_{1})}{\sum_{\tilde{x}_{1}}\cancel{\big{[}\prod_{i\in M_{t}}(1-t)\prod_{i\in U_{t}}t\big{]}}\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=\tilde{x}_{1}^{i})\big{]}p(t)p(\tilde{x}_{1})}$ (9)
$=\frac{\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=x_{1}^{i})\big{]}p(x_{1})}{\sum_{\tilde{x}_{1}}\big{[}\prod_{i\in U_{t}}\mathbb{I}(x_{t}^{i}=\tilde{x}_{1}^{i})\big{]}p(\tilde{x}_{1})}$ (10)
$=p(x_{1}\mid x_{t}),$ (11)
where $\mathbb{I}$ is the indicator function. Thus, we can rewrite the masking flow matching loss function as:
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t},t)\right]$ (12)
$=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p^{\theta}(x_{1}^{i}|x_{t})\right]$ (13)
@@
Equivalence between flow matching on discrete state-space and order-agnostic autoregressive models
Order-agnostic autoregressive models (OA-AR) differ from standard autoregressive models by training on arbitrary permutations of the sequence ordering. We will now demonstrate that under specific conditions, the discrete flow matching objective is equivalent to the order-agnostic autoregressive objective.
@@
$p^{\theta}(x_{1})=\prod_{i=1}^{D}p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})$ (15)
The OA-AR objective is to maximize the expected log-likelihood over all possible permutations:
$\mathcal{L}_{\text{OA-AR}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\sum_{i=1}^{D}\log p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})\right]$ (16)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D)),i\sim\text{Unif}(\{1,2,\ldots,D\})}\left[\log p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})\right]$ (17)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),i\sim\text{Unif}(\{1,2,\ldots,D\}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\log p^{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})\right]$ (18)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim\text{Unif}([0,1]),x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p^{\theta}(x_{1}\mid x_{t})\right]$ (19)
@@
Relationship to autoregressive models and discrete time discrete state-space diffusion models
@@
Autoregressive (AR) models represent another important class of generative models for discrete spaces that can be connected to the diffusion framework. AR models factorize the joint distribution sequentially as $p(x_{1})=\prod_{i=1}^{D}p(x_{1}^{i}|x_{1}^{1},x_{1}^{2},...,x_{1}^{i-1})$ , generating tokens one at a time in a fixed order. Austin et al. [17] showed that AR models can be interpreted as discrete space diffusion models with a particular form of noising process.
@@
For AR models, Yang and Klein [22] introduced FUDGE, a method that performs guidance by modulating each conditional distribution using a learned classifier evaluated on the partially completed sequence. FUDGE can be viewed as a special case of ProteinGuidewhen applied specifically to AR models. While discrete time diffusion models are trained with a different objective than their continuous-time counterparts, they can still benefit from our guidance methodology. The learned denoising model from a discrete-time training objective can be used to induce a rate matrix for sampling by assuming smooth interpolation between discrete time points. However, both theoretical analysis and empirical results [32, 34, 38, 39] suggest that models trained with continuous-time objectives often achieve superior performance in practice.
@@
This relationship between different model classes highlights the generality of ProteinGuide as a framework that can accommodate various generative architectures, including AR models, discrete-time diffusion models, and the continuous-time formulations we primarily focus on in this work.
@@
Guidance is a useful technique that enables conditional generation with diffusion and flow models. Originally developed for continuous state-space models, guidance allows steering the generative process toward samples with desired properties without requiring conditional retraining [13, 15, 14]. In continuous state-space diffusion models, the generation process involves sampling from a reverse stochastic differential equation (SDE) that gradually transforms noise into structured data. To condition this process on a desired property $y$, it has been shown that one can leverage Bayes’ theorem to obtain the conditional score function:
@@
This formulation, termed classifier guidance, combines the unconditional score $\nabla_{x_{t}}\log p_{t}(x_{t})$ with the score of a predictor model $\nabla_{x_{t}}\log p_{t}(y|x_{t})$. The strength of guidance can be controlled by introducing a parameter $\gamma$:
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\gamma\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (22)
@@
While guidance, either classifier-based, classifier-free or autoguidance, has proven extremely valuable for continuous state-space models, extending these concepts to discrete state-spaces poses unique challenges, since the score functions are not defined and these formulations cannot be directly applied. We sidestep the need for a score function by instead returning to Bayes’ theorem, which, in the context of diffusion on discrete state-spaces, dictates that we must effectively compute
@@
[…]
To perform predictor guidance in this setting, we derive a principled approach to adjust these rates based on a predictor model $p(y|x_{t})$ that evaluates how likely a state $x_{t}$ is to have the desired property $y$. Specifically, we show that the conditional rates can be obtained as:
[…]
for any state $\tilde{x}\neq x$. This formulation directly implements Bayes’ theorem at the level of transition rates, modulating transitions toward states that have higher likelihood of possessing the desired property.
Similar to continuous state-space guidance, we can control the strength of guidance with a parameter $\gamma$:
[…]
, or derive predictor-free guidance discrete state-spaces:
@@
For computational efficiency, especially with large state-spaces, we introduced Taylor-approximated guidance (TAG), which approximates the likelihood ratio using a first-order Taylor expansion inspired by Grathwohl et al. [44]:
@@
We demonstrated how assay-labeled experimental data could be used to condition generative models for protein sequences, improving the success rates for such models in generating sequences with the desired function. To illustrate this use case, we guided ProteinMPNN with a large dataset of experimental folding stability measurements on mini-proteins [23] to generate sequences with improved stability. Stability is reported in real-valued units of $\Delta\Delta G$, which quantifies the difference in stability from the wild-type sequence. A protein is considered stable if it is at least as stable as the wild-type (i.e., $\Delta\Delta G\geq 0$). To generate sequences with improved stability, we followed the experimental setup from Nisonoff et al. [16]. Specifically, we use ProteinMPNN [1] as the inverse folding model, $p(\mathbf{s}\,|\,\mathbf{x})$, where $\mathbf{s}$ corresponds to a protein sequence and $\mathbf{x}$ is the desired backbone structure. Then we apply guidance with the stability predictor from Nisonoff et al. [16], $p(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$, to obtain a stability-conditioned inverse folding model, $p(\mathbf{s}\,|\,\mathbf{x},\Delta\Delta G\geq 0)$.

### v2

@@ Methods
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right].$ (7)
Our goal is to demonstrate that this loss function is equivalent to the loss function used by a class of Masked Language Models (MLMs), that we will refer to as “generative Masked Language Models”. The key distinction between generative Masked Language Models and their non-generative counterparts, is that during training, generative MLMs mask tokens with a masking probability, $t\in[0,1]$, sampled from a distribution $p(t)$ that has full support on the interval $[0,1]$. This requirement is met by MLMs such as ESM3 [4], but, importantly, is not met by BERT-style MLMs such as ESM2 [44], since these models use a fixed percentage of tokens that are masked at each step. Letting $x_{t}$ refer to the masked sequence, we have that the generative masked language modeling objective function is
$\mathcal{L}_{\text{MLM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t})\right],$ (8)
@@
For any position $i\in U_{t}$, we know with certainty that $x_{1}^{i}=x_{t}^{i}$ by the definition of our forward process. The key observation is that the time variable $t$ only influences which tokens are masked (the masking pattern), but once we observe $x_{t}$, we have complete information about which positions are masked and which are not. The time variable, therefore, provides no additional information needed to predict the clean data. Using a mathematical argument first proposed by Gat et al. [38], we can prove this as follows:
@@
[…]
$=\frac{\big{[}\prod_{i\in U_{t}}\delta\{x_{t}^{i},x_{1}^{i}\}\big{]}p(x_{1})}{\sum_{\tilde{x}_{1}}\big{[}\prod_{i\in U_{t}}\delta\{x_{t}^{i},\tilde{x}_{1}^{i}\}\big{]}p(\tilde{x}_{1})}$ (12)
$=p(x_{1}\mid x_{t}).$ (13)
where Equation 11 uses the factorization over the dimensions and analyze separately the two cases of whether a position is masked or unmasked, and Equation 12 shows that $p(x_{1}\,|\,x_{t},t)$ does not depend on $t$, giving us Equation 13. Thus, we can rewrite the masking flow matching loss function as:
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right]$ (14)
$=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t})\right]$ (15)
@@
Equivalence between flow matching on discrete state-spaces and order-agnostic autoregressive models
Order-agnostic autoregressive models (OA-AR) differ from standard autoregressive models by training on arbitrary permutations of the sequence ordering [42, 20]. We will now demonstrate that under a masking noise process, the discrete flow matching objective is equivalent to the order-agnostic autoregressive objective.
@@
$p_{\theta}(x_{1})=\prod_{i=1}^{D}p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})$ (17)
where $x_{1}^{\sigma(<i)}$ denotes $(x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})$. The OA-AR objective is to maximize the expected log-likelihood over all possible permutations:
$\mathcal{L}_{\text{OA-AR}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\sum_{i=1}^{D}\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (18)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D)),i\sim\text{Unif}(\{1,...,D\})}\left[\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (19)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),i\sim\text{Unif}(\{1,...,D\}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (20)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim\text{Unif}([0,1]),x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p_{\theta}(x_{1}\mid x_{t})\right]$ (21)
@@
Decoding ordering in order-agnostic autoregressive models
While the mathematical equivalence between order-agnostic autoregressive (OA-AR) models and discrete flow matching is exact in theory, a practical subtlety arises when implementing OA-AR models with causal attention mechanisms. In practice, many OA-AR models, including ProteinMPNN [1], are implemented using standard autoregressive architectures with causal attention masks. When computing the probability $p_{\theta}(x_{1}^{i}|x_{t})$ for a masked position $i$ given a partially masked sequence $x_{t}$, the model must process the unmasked positions in $x_{t}$ in some specific order due to the causal nature of the attention mechanism. Formally, for a position $i\in M_{t}$ (the set of masked positions), the probability $p_{\theta}(x_{1}^{i}|x_{t})$ should theoretically be independent of how the unmasked positions in $U_{t}={j:x_{t}^{j}\neq\text{?}}$ are ordered. However, in practice, the causal attention mechanism means that:
$p_{\theta}(x_{1}^{i}|x_{t})=p_{\theta}(x_{1}^{i}|x_{t}^{\sigma(1)},x_{t}^{\sigma(2)},\ldots,x_{t}^{\sigma(|U_{t}|)})$ (23)
where $\sigma$ is some permutation of the unmasked positions in $U_{t}$, and the probability depends on this specific ordering $\sigma$. During training, the OA-AR objective encourages the model to produce consistent predictions regardless of the permutation $\sigma$. The expectation over all possible permutations in the training objective should, in principle, ensure that $p_{\theta}(x_{1}^{i}|x_{t})$ is invariant to the ordering of unmasked positions. However, in practice, this invariance may not be perfectly achieved due to imperfect optimization and finite-sample training. To address this subtlety and ensure consistent behavior during guided sampling, for the protein stability experiment that guides ProteinMPNN, we fix the decoding ordering for $x_{t}$ based on the residue positions of the unmasked tokens. Specifically, we sort the unmasked positions in $U_{t}$ in ascending order of their sequence positions:
$\sigma(j)=\text{sort}(U_{t})[j]$ (24)
where $\text{sort}(U_{t})$ returns the unmasked positions ordered by their indices in the original sequence. This approach provides a deterministic and reproducible ordering for any given masked state $x_{t}$, respects the natural left-to-right order of residues in the protein sequence, and requires no additional hyperparameters or complex ordering strategies. By using this fixed position-based ordering, we ensure that the conditional probabilities $p_{\theta}(x_{1}^{i}|x_{t})$ are well-defined and consistent across different sampling runs, enabling reliable application of the guidance method to OA-AR models trained with causal attention.
@@
Relationship to autoregressive models and discrete time diffusion models on discrete state-spaces
@@
Autoregressive (AR) models represent another important class of generative models for discrete state-spaces that can be connected to the diffusion framework. AR models factorize the joint distribution sequentially as $p(x_{1})=\prod_{i=1}^{D}p(x_{1}^{i}|x_{1}^{1},x_{1}^{2},...,x_{1}^{i-1})$ , generating tokens one at a time in a fixed order. Austin et al. [19] showed that AR models can be interpreted as discrete space diffusion models with a particular form of noising process. For AR models, Yang and Klein [24] introduced FUDGE, a method that performs guidance by modulating each conditional distribution using a learned classifier evaluated on the partially completed sequence. FUDGE can be viewed as a special case of ProteinGuide when applied specifically to AR models.
@@
While discrete time diffusion models [19] are trained with a different objective than their continuous-time counterparts, they can still benefit from our guidance methodology. The learned denoising model from a discrete-time training objective can be used to induce a rate matrix for sampling by assuming smooth interpolation between discrete time points. However, both theoretical analysis and empirical results [31, 33, 37, 38] suggest that models trained with continuous-time objectives often achieve superior performance in practice.
@@
This relationship between different model classes highlights the generality of ProteinGuide as a framework that can accommodate various generative architectures, including AR models, discrete-time diffusion models in discrete state-spaces, and the continuous-time formulations we primarily focus on in this work.
@@
Guidance is a useful technique that enables conditional generation with diffusion and flow models [13, 15, 14]. Originally developed for continuous state-space models, guidance allows steering the generative process toward samples with desired properties without requiring conditional training. In continuous state-space diffusion models, the generation process involves sampling from a reverse stochastic differential equation (SDE) that gradually transforms noise into data [15]. To condition this process on a desired property $y$, it has been shown that one can leverage Bayes’ theorem to obtain the conditional score function [15]:
@@
This formulation, termed classifier guidance, combines the unconditional score $\nabla_{x_{t}}\log p_{t}(x_{t})$ with the gradient of a predictor model $\nabla_{x_{t}}\log p_{t}(y|x_{t})$. The strength of guidance can be controlled by introducing a guidance strength parameter $\gamma$ [14]:
$\nabla_{x_{t}}\log p^{(\gamma)}_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\gamma\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (26)
@@
While guidance, either classifier-based, classifier-free or autoguidance, has proven extremely valuable for continuous state-space models, extending these concepts to discrete state-spaces poses unique challenges, since the score functions are not defined and these formulations cannot be directly applied. One can sidestep the need for a score function by instead returning to Bayes’ theorem, which, in the context of diffusion on discrete state-spaces, dictates that we must effectively compute
@@
Here we provide a brief overview of the results presented in Nisonoff et al. [16]. To enable guidance for discrete state-space models, we turn to continuous-time diffusion models and flow models in discrete state-spaces. As described above, these models are based on continuous-time Markov chains (CTMCs) and therefore parameterize transition rates $R_{t}(x,\tilde{x})$ rather than score functions. These models also assume that the noising processes in each dimension are independent of one another (but not necessarily identical), which has the important consequence that the probability that two or more dimensions of the current state jump (transition) at exactly the same time is zero, which makes these models tractable [31, 37]. It turns out that the same assumption also makes guidance tractable for these models [16].
To perform predictor guidance in this setting, we derived a principled approach to adjust these rates based on a predictor model $p(y|x_{t})$ that evaluates how likely a state $x_{t}$ is to have the desired property $y$ [16]. Specifically, we showed that the conditional rates can be obtained as:
$R_{t}(x,\tilde{x}|y)=\frac{p(y|\tilde{x},t)}{p(y|x,t)}R_{t}(x,\tilde{x})$ (28)
for any state $\tilde{x}$ that differs from $x$ in at most one dimension. This formulation directly implements Bayes’ theorem at the level of transition rates, modulating transitions toward states that have higher likelihood of possessing the desired property.
Similar to continuous state-space guidance, we can control the strength of guidance with a guidance strength parameter $\gamma$:
$R_{t}^{(\gamma)}(x,\tilde{x}|y)=\left[\frac{p(y|\tilde{x},t)}{p(y|x,t)}\right]^{\gamma}R_{t}(x,\tilde{x})$ (29)
or derive predictor-free guidance discrete state-spaces:
@@
For computational efficiency, especially with large state-spaces, we introduced Taylor-approximated guidance (TAG) [16], which approximates the likelihood ratio using a first-order Taylor expansion inspired by Grathwohl et al. [47]:
@@
Additional discussion of related work
Several recent works have proposed related but different techniques from Nisonoff et al. [16] for steering discrete state-space diffusion models. These approaches can be broadly categorized into guidance-based methods (which modify the sampling process at inference time) and fine-tuning approaches (which update the parameters of the pre-trained model).
Li et al. [49] proposed a derivative-free sampling technique for reward maximization in both continuous and discrete spaces using Sequential Monte Carlo (SMC). Their approach uses the unconditional model as the proposal distribution and reweights multiple particles with the reward function. Singhal et al. [51] built upon and generalized the methods in Wu et al. [48] and Li et al. [49] to enable more flexible choices over the sampling hyperparameters. Schiff et al. [50] presented a discrete-time analog of the continuous-time guidance method in Nisonoff et al. [16]. As noted in Schiff et al. [50], in discrete time, the normalizing constant in Equation 27 is still intractable, necessitating additional assumptions to achieve tractability, whereas Nisonoff et al. [16] showed that the continuous-time formulation of discrete diffusion models provides exactly the tractability necessary to perform exact guidance. In addition, the continuous-time formulation naturally enables guidance for discrete state-space flow matching models, whereas the discrete-time formation does not. Lee et al. [52] observed that when applying stronger guidance, sampling can be improved with SMC, aligning with challenges in low-temperature sampling noted in Du et al. [53] and Ingraham et al. [3]. For AR models, Zhao et al. [54] proposed an SMC-based steering method that learns intermediate twist functions via contrastive learning, similar to learning time-dependent predictors in diffusion guidance.
Recent works have also explored updating model parameters to incorporate conditioning rather than modifying the sampling process. Wang et al. [55] proposed to use the Gumbel-Softmax trick to make discrete diffusion trajectories differentiable and enable backpropagation of rewards through entire sampling paths. Concurrently, Rector-Brooks et al. [56] proposed to frame the steering problem as sampling from a Bayesian posterior where the pre-trained model serves as the prior and the reward acts as the likelihood. Both fine-tuning approaches established connections to Nisonoff et al. [16] while offering complementary perspectives on solving the steering problem.
Direct Preference Optimization (DPO [57]) and related work [11, 58] enable updating of pre-trained models, although up until recently, these methods have been applicable only to models with tractable likelihoods, thereby restricting them largely to autoregressive models. For example, Widatalla et al. [11] applied DPO to ESM-IF1, an autoregressive inverse folding model, to incorporate experimental stability data. In another example, Stocco et al. [59] applied DPO to an autoregressive model conditionally trained on enzyme classes to guide the generation towards various enzyme properties. Wallace et al. [60] extended DPO to diffusion models on continuous state-spaces by way of approximation, while Borso et al. [61] explored how to similarly do so for these models on discrete state-spaces. Importantly, the connections previously mentioned between MLMs, diffusion and flow matching models on discrete state-spaces, and order-agnostic autoregressive models, means that we can also generalize the aforementioned approximate DPO to these same model classes. Herein, we focus our study on plug-and-play guidance, rather than DPO-style approaches, although the latter could prove equally useful. Recent work by Yang et al. [62] compared various guidance-based and fine-tuning based approaches on several in-silico tasks for protein fitness optimization and found guidance-based approaches for discrete state-space diffusion models to be generally more effective than fine-tuning based approaches. In addition, they found that guidance-based approaches require less hyper-parameter tuning and have lower computational costs, making them more practically accessible to everyday users.
Reinforcement Learning from Human Feedback (RLHF) offers an alternative approach to model alignment through methods like Proximal Policy Optimization (PPO) [63]. PPO does not technically require the retrained model to have tractable likelihoods, but it requires an explicit reward model and on-policy training, making it often more complicated to train than more direct approaches like DPO [57]. In addition, the standard RLHF pipeline typically begins with more heuristic methods such as continuing to train an existing model with new data (i.e., supervised fine-tuning), for which the statistical outcome is not well-defined, and for which important information from the already-trained model may be forgotten. Blalock et al. [64] recently applied PPO to align protein language models e.g., ESM2 with experimental measurements. Notably, this approach does not directly sample from generative models in the conventional sense; instead, it employs a position-wise scoring strategy where models evaluate mutation probabilities at individual positions sequentially, followed by an iterative selection process that combines high-scoring mutations.
There is another line of work that model discrete data by embedding the data into continuous state-spaces, either on a probabilistic simplex [65] or on a continuous latent space [66, 67]. Some of the approaches proposed to apply guidance to diffusion models on the continuous state-space representations of discrete state-space objects in order to utilize the guidance approaches for continuous state-space diffusion. For example, Gruver et al. [68] propose to perform guidance in the hidden layers of the unconditional diffusion model by jointly training the unconditional denoising model and the classifier on one hidden layer of the denoising model. In another example, Lisanza et al. [69] propose to fine-tune RoseTTAFold with gaussian diffusion on the one-hot encoded sequence space in order to perform guidance in this continuous state-space. We note that approaches which perform guidance in continuous spaces lose the discrete structure of the data during generation, which can be important for settings where the discrete structures contain information that are useful for guidance [37].
@@
We demonstrated how assay-labeled experimental data could be used to condition generative models for protein sequences, improving the success rates for such models in generating sequences with the desired function. To illustrate this use case, we guided ProteinMPNN with a large dataset of experimental folding stability measurements on mini-proteins [25] to generate sequences with improved stability. Stability is reported in real-valued units of $\Delta\Delta G$, which quantifies the difference in stability from the wild-type sequence. A protein is considered stable if it is at least as stable as the wild-type (i.e., $\Delta\Delta G\leq 0$). To generate sequences with improved stability, we followed the experimental setup from Nisonoff et al. [16]. Specifically, we used ProteinMPNN [1] as the inverse folding model, $p_{\theta}(\mathbf{s}\,|\,\mathbf{x})$, where $\mathbf{s}$ corresponds to a protein sequence and $\mathbf{x}$ is the desired backbone structure. As described above, the OA-AR nature of ProteinMPNN means we can treat the pre-trained model as a flow matching / diffusion model, $p_{\theta}(\mathbf{s}_{1}\,|\,\mathbf{s}_{t},\mathbf{x})$, under a masking noise process. Then we apply guidance with the stability predictor from Nisonoff et al. [16], $p_{\phi}(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$, to obtain a stability-conditioned inverse folding model, $p_{\theta,\phi}(\mathbf{s}\,|\,\mathbf{x},\Delta\Delta G\leq 0)$.